Circular No. 6967

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980703 M. R. Zapatero Osorio, Instituto de Astrofísica de Canarias (IAC);

A. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofísica Espacial y Física Fundamental; A. Oscoz and S. Kemp, IAC; and F. Frontera and L. Nicastro, on behalf of the BeppoSAX team, report: “Images in the R band have been obtained at the 0.8-m IAC80 telescope at Teide Observatory on July 4 and 6 of the field including the BeppoSAX error box for the x-ray source 1SAX J2359.1+0835 (cf. GCN 127) that is presumably related to GRB 980703, which was detected by both BATSE and RXTE. These data reveal an object located at α = 23^{h}59^{m}06^{s}.7, δ = +8^{o}35′07′′ (equinox 2000.0; ± 2′′), which has faded by ∼ 1.5 mag at this wavelength. A preliminary photometric analysis yields R = 20.6 ± 0.1 on July 4.123 UT and 22.1 ± 0.2 on July 6.137. No other objects varying by > 0.2 mag are seen down to a limiting magnitude of R = 22.5. A finder chart is available via anonymous

ftp to ftp.ll.iac.es (cd publong/grb980703/).”

CH CYGNI M. Mikolajewski, Center for Astronomy, Torun; and L. Leedjarv, Tartu

Observatory, communicate: “A new outburst of the symbiotic star CH Cyg seems to have begun. Photoelectric UBVRI (Johnson) observations on the Torun observatory 0.6-m telescope show significant development of the hot continuum since mid-May: May 15.87 UT, U = 10.84, B = 10.46, V = 8.69, R = 5.04, I = 2.40; 19.99, 10.58, 10.39, 8.67, 5.09, 2.42; June 4.02, 9.91, 10.03, 8.54, 5.20, 2.63; 27.91, 9.22, 9.58, 8.42, 5.31, 2.82; July 3.01, 8.95, 9.32, 8.23, 5.33, 2.87. During three nights (June 21, 27, 28), ∼ 3 hr of observing on each night have revealed flickering on a timescale of 5–15 min, with amplitudes of ∼ 0.50–0.40, 0.40–0.30, 0.25–0.20, and 0.08– 0.06 magnitude in U, B, V, and R, respectively. Spectra obtained on the 1.5-m telescope of Tartu Observatory on June 27.91 show remarkable strengthening of the emission lines. Hα and other Balmer lines show a double-peaked profile, with a separation of peaks of ∼ 100 km/s and with the violet peak stronger than the red one. Strong emission lines of He I (388.9, 402.6, 447.1 nm) and Fe II have become visible. Forbidden emission lines of [Ne III] (386.8 nm), [S II] (406.8 nm), [O III] (500.7 nm), and [Fe II] (many) have remained strong, while those of [N II] (654.8 and 658.3 nm) have become considerably weaker.”

1998 July 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green